OpenStudy (anonymous):

49^(x+1)=square root of 7?

OpenStudy (ness9630):

Get common bases, make both have a base of 7.

OpenStudy (campbell_st):

well rewrite each side of the equation in index form with the same base \[(7^2)^{(x + 1)} = 7^{\frac{1}{2}}\] which simplifies to \[7^{2x + 2} = 7 ^{\frac{1}{2}}\] now just equate the powers and solve for x

OpenStudy (anonymous):

its \[\sqrt{7}\] not 7 \[\frac{ 1 }{ 2 } \]

OpenStudy (campbell_st):

well well a little index information \[\sqrt{7} = 7^{\frac{1}{2}}\] to test it get a calculator and find \[\sqrt{7} =\] next, using the power key on your calculator find \[7^{\frac{1}{2}}.... or.... 7^{0.5}\] what happens...

OpenStudy (anonymous):

oh!! okay so now ive got 2x+2 =\[\frac{ 1 }{ 2 }\] now what?

OpenStudy (campbell_st):

solve for x

OpenStudy (akashdeepdeb):

@shaa Need help?

OpenStudy (anonymous):

I got -1.25! that's wrong!

OpenStudy (anonymous):

yes please explain!

OpenStudy (akashdeepdeb):

Hold on for a sec... The answer should be.....

OpenStudy (akashdeepdeb):

Is there an option such as -3/4?

OpenStudy (anonymous):

Ive got no options

OpenStudy (campbell_st):

well the answer may need to be in fraction form to start so 2x + 2 = 1/2 subtract 2 2x = -3/2

OpenStudy (akashdeepdeb):

-0.75 ?

OpenStudy (akashdeepdeb):

Okay follow him. :D

OpenStudy (anonymous):

so the answer is -1.75?

OpenStudy (campbell_st):

no... you have 2x = -3/2 divide both sides of the equation by 2 x = -3/4 to check if its correct, substitute into the original equation...

OpenStudy (anonymous):

-0.75 I mean

OpenStudy (anonymous):

sorry I typed it wrong haha! its -0.75 right?

OpenStudy (campbell_st):

well i'd probably write the answer is -3/4... but otherwise you're correct

OpenStudy (anonymous):

finally! thanks so much for your help! :)

OpenStudy (campbell_st):

good luck with your maths